# Panning for gold with the Neil Gehrels Swift Observatory: an optimal strategy for finding the counterparts to gravitational wave events

### v1

@@ Abstract
The LIGO, Virgo and KAGRA gravitational wave observatories are currently undertaking their O4 observing run offering the opportunity to discover new electromagnetic counterparts to gravitational wave events. We examine the capability of the Neil Gehrels Swift Observatory (Swift) to respond to these triggers, primarily binary neutron star mergers, with both the UV/Optical Telescope (UVOT) and the X-ray Telescope (XRT). We simulate Swift’s response to a trigger under different strategies using model skymaps, convolving these with the 2MPZ catalogue to produce an ordered list of observing fields, deriving the time taken for Swift to reach the correct field and simulating the instrumental responses to modelled kilonovae and short gamma-ray burst afterglows. We find that UVOT, using the u filter with an exposure time of order 120 s, is optimal for most follow-up observations and that we are likely to detect counterparts in $\sim 6$% of all binary neutron star triggers. We find that the gravitational wave 90% error area and measured distance to the trigger allow us to select optimal triggers to follow-up. Focussing on sources less than 300 Mpc away or 500 Mpc if the error area is less than a few hundred square degrees, distances greater than previously assumed, offer the best opportunity for discovery by Swift with $\sim 5-30$% of triggers having detection probabilities $\geq 0.5$. At even greater distances, we can further optimise our follow-up by adopting a longer 250 s or 500 s exposure time.
@@ 1 Introduction
GW 170817 was a true watershed moment in astronomy, the first time a gravitational wave (GW) detection was found to have an electromagnetic (EM) counterpart (e.g. Abbott et al., 2017; Covino et al., 2017; Cowperthwaite et al., 2017; Evans et al., 2017; Goldstein et al., 2017; Kilpatrick et al., 2017; Savchenko et al., 2017; Smartt et al., 2017; Tanvir et al., 2017). However, at time of writing, it remains a unique event. The Laser Interferometer Gravitational-Wave Observatory (LIGO), the Virgo interferometer and Kamioka Gravitational Wave Detector (KAGRA), together referred to in this work as LVK, commenced their O4 run on 24 May 2023 offering the opportunity to finally make GW 170817 commonplace rather than unique. It is therefore crucial to devise an optimal strategy to follow-up GW detections and quickly discover any EM counterparts. Here we build on the work of Evans et al. (2016a, hereafter Paper I) and Evans et al. (2016b, hereafter Paper II, see also ), which discuss the strategy for previous LVK runs. However, both works assumed that the afterglow from a short GRB would be the signal most likely to be detected by Swift and therefore concentrated on Swift’s X-Ray Telescope (XRT) as the discovery instrument. However, the afterglow of GW 170817 was not detected by the XRT [^1] Although we note that Swift didn’t observe its position until $\sim 14$ hours after the trigger. but did have a UV bright kilonova (Evans et al., 2017) detected by the UV/Optical Telescope (UVOT).
@@
A short GRB and a kilonova are the two EM transients that are typically expected to accompany the merger of a binary neutron star (BNS) or neutron star-black hole (NSBH) system, in addition to the gravitational waves detectable by LVK. Short GRBs result in two signals, the prompt gamma-ray emission that emerges from within a relativistic jet (the burst itself) and a synchrotron powered broadband afterglow resulting from jet interactions with the circumburst medium. However due to Doppler beaming, both these components are strongly directional, particularly at the wavelengths accessible to Swift. The narrow opening angle of the jet in a short GRB ($\sim 15\degr$, Fong et al., 2015) therefore means that only a very small fraction of the merger population detectable by LVK through relatively isotropic gravitational wave emission will have bright and detectable afterglows. Kilonovae, on the other hand, are powered by r-process nucleosynthesis in the merger ejecta and emit far more isotropically (e.g. Metzger, 2019). While there is still significant viewing angle dependence due to the nature of the ejecta, a search focussed on kilonova emission might therefore be most effective. In this work, therefore, we focus primarily on the UVOT but also apply our analysis to the XRT.
@@
In this paper, we examine the UVOT’s capability in detecting a counterpart from a GW trigger during O4, specifically a BNS merger and the resultant kilonova. Although we generally find the probability of a significant detection relatively low, we can improve our chances by using specific criteria such as the 90% error area and the luminosity distance. This was already used to inform Swift’s strategy and this work has aided in updating these criteria and the strategy accordingly. Selecting events using these criteria significantly enhances the chance of UVOT promptly discovering an EM counterpart and ensures Swift remains effective across all aspects of its mission.
@@ 2.2 Host selection
For a given trigger, therefore, the probability of it being in a known galaxy is simply the completeness at that distance i.e. $P_{\rm gal}=C$. If probabilistically determined to be located in a known galaxy [^4] Due to high natal kick velocities, a trigger could occur a significant distance ‘outside’ its host galaxy (e.g. Mandhai et al., 2022). However, at the luminosity distances involved, this offset is relatively insignificant and can therefore be neglected in our analysis., a catalogued galaxy is randomly selected to be the host weighted by the K band luminosity as a proxy for the galaxies’ stellar masses. We also require the host to be at an appropriate distance and a similar declination to the original position of the trigger. The original skymap is then shifted across the sky to match the trigger’s new location (as the real position of the source is known relative to the skymap). For sources that weren’t selected to be in a host galaxy, their position and associated skymap were left unchanged from the original simulation.
@@ 3.1 Kilonova component
Their tiny observed population mean that kilonovae remain poorly understood and models of their emission are diverse in behaviour (e.g. Rossi et al., 2022). For instance, toy models with assumed ejecta velocity distributions and opacities can approximate much of the behaviour of a kilonova (Metzger, 2019). More advanced radiation transfer models, on the other hand, can fully predict the angular dependence of luminosity and spectra (Kasen et al., 2017; Bulla, 2019, 2023). Arguably the most successful models, however, combine inferences and predictions from both the GW and EM data (Coughlin et al., 2019; Dietrich et al., 2020; Breschi et al., 2021). Here we use the analytical BNS models of Nicholl et al. (2021, hereafter N21) which offer a distinct advantage over these options - they are designed to generate a light curve from only parameters inferred from a GW signal. These models are therefore ideal for use in our planning and during O4 itself.
@@ 5.2 Binary black hole mergers
To this point we have only considered the gravitational wave sources detectable by LVK that are expected to produce significant EM emission: BNS and NSBH mergers. However, the vast majority of sources detected by LVK are binary black hole (BBH) mergers. Rather than the kilonova and short GRB afterglow emission components examined here, the possible EM emission from a BBH merger is far more unconstrained. Due to the inherent nature of a BBH system, EM emission can only be produced if the merger occurs in a dense environment, for instance, the disk of an active galactic nucleus (e.g. Kelly et al., 2017; McKernan et al., 2019). No confirmed EM counterpart to a BBH merger has thus far been discovered with only a few examples of possible candidates (e.g. Connaughton et al., 2016; Graham et al., 2020, 2023, see also Greiner et al. 2016 and Connaughton et al. 2018). Due to the lack of constraints on these triggers’ emission, Swift follows a different strategy to that explored in this work. Instead Swift follows-up only the best localised events with 90% error areas $\leq 30$ deg2and using minimal length 80 s exposures. This ensures the area is covered rapidly and minimises disruption to Swift’s other activities. During O4, Swift has followed up several BBH mergers meeting the criterion above, with no likely EM counterpart candidates so far.
@@ Acknowledgements
RAJEF acknowledges support from the UK Space Agency and the European Union’s Horizon 2020 Programme under the AHEAD2020 project (grant agreement number 871158).

### v2

@@ Abstract
The LIGO, Virgo and KAGRA gravitational wave observatories are currently undertaking their O4 observing run offering the opportunity to discover new electromagnetic counterparts to gravitational wave events. We examine the capability of the Neil Gehrels Swift Observatory (Swift) to respond to these triggers, primarily binary neutron star mergers, with both the UV/Optical Telescope (UVOT) and the X-ray Telescope (XRT). We simulate Swift’s response to a trigger under different strategies using model skymaps, convolving these with the 2MPZ catalogue to produce an ordered list of observing fields, deriving the time taken for Swift to reach the correct field and simulating the instrumental responses to modelled kilonovae and short gamma-ray burst afterglows. We find that UVOT, using the u filter with an exposure time of order 120 s, is optimal for most follow-up observations and that we are likely to detect counterparts in $\sim 6$% of all binary neutron star triggers detectable by LVK in O4. We find that the gravitational wave 90% error area and measured distance to the trigger allow us to select optimal triggers to follow-up. Focussing on sources less than 300 Mpc away, or 500 Mpc if the error area is less than a few hundred square degrees, distances greater than previously assumed, offer the best opportunity for discovery by Swift with $\sim 5-30$% of triggers having detection probabilities $\geq 0.5$. At even greater distances, we can further optimise our follow-up by adopting a longer 250 s or 500 s exposure time.
@@ 1 Introduction
GW 170817 was a true watershed moment in astronomy, the first time a gravitational wave (GW) detection was found to have an electromagnetic (EM) counterpart (e.g. Abbott et al., 2017; Covino et al., 2017; Cowperthwaite et al., 2017; Evans et al., 2017; Goldstein et al., 2017; Kilpatrick et al., 2017; Savchenko et al., 2017; Smartt et al., 2017; Tanvir et al., 2017; Troja et al., 2017). However, at time of writing, it remains a unique event. The Laser Interferometer Gravitational-Wave Observatory (LIGO), the Virgo interferometer and Kamioka Gravitational Wave Detector (KAGRA), together referred to in this work as LVK, commenced their O4 run on 24 May 2023 offering the opportunity to finally make GW 170817 commonplace rather than unique. It is therefore crucial to devise an optimal strategy to follow-up GW detections and quickly discover any EM counterparts. Here we build on the work of Evans et al. (2016a, hereafter Paper I) and Evans et al. (2016b, hereafter Paper II, see also ), which discuss the strategy for previous LVK runs. However, both works assumed that the afterglow from a short GRB would be the signal most likely to be detected by Swift and therefore concentrated on Swift’s X-Ray Telescope (XRT) as the discovery instrument. However, the afterglow of GW 170817 was not detected by the XRT [^1] Although we note that Swift didn’t observe its position until $\sim 14$ hours after the trigger. but did have a UV bright kilonova (Evans et al., 2017) detected by the UV/Optical Telescope (UVOT).
@@
A short GRB and a kilonova are the two EM transients that are typically expected to accompany the merger of a binary neutron star (BNS) or neutron star-black hole (NSBH) system, in addition to the gravitational waves detectable by LVK. Short GRBs result in two signals, the prompt gamma-ray emission that emerges from within a relativistic jet (the burst itself) and a synchrotron powered broadband afterglow resulting from jet interactions with the circumburst medium. However due to Doppler beaming, both these components are, at least initially, strongly directional, particularly at the wavelengths accessible to Swift. While structured jets or cocoon emission (Nakar & Piran, 2017; Piro & Kollmeier, 2018) do result in more isotropic emission, the relatively narrow opening angle of the jet in a short GRB ($\sim 15\degr$, Fong et al., 2015) therefore means that only a very small fraction of the merger population detectable by LVK through the far more isotropic gravitational wave emission will have bright and detectable afterglows. Kilonovae, on the other hand, are powered by r-process nucleosynthesis in the merger ejecta and emit much more isotropically (e.g. Metzger, 2019). While there is still significant viewing angle dependence due to the nature of the ejecta, a search focussed on kilonova emission might therefore be most effective. In this work, therefore, we focus primarily on the UVOT but also apply our analysis to the XRT.
@@
In this paper, we examine the UVOT’s capability in detecting a counterpart from a GW trigger during O4. We concentrate primarily on BNS mergers as their resultant kilonovae are significantly brighter and easier to detect than those expected from NSBH mergers (e.g. Gompertz et al., 2023), although we also briefly discuss these sources. Although we generally find the probability of a significant detection relatively low, we can improve our chances by using specific criteria such as the 90% error area and the luminosity distance. This was already used to inform Swift’s strategy and this work has aided in updating these criteria and the strategy accordingly. Selecting events using these criteria significantly enhances the chance of UVOT promptly discovering an EM counterpart and ensures Swift remains effective across all aspects of its mission.
@@ 2.2 Host selection
For a given trigger, therefore, the probability of it being in a known galaxy is simply the completeness at that distance i.e. $P_{\rm gal}=C$. If probabilistically determined to be located in a known galaxy [^4] Due to high natal kick velocities, a trigger could occur a significant distance ‘outside’ its host galaxy (e.g. Mandhai et al., 2022). However, at the luminosity distances involved, this offset is relatively insignificant and can therefore be neglected in our analysis., a catalogued galaxy is randomly selected to be the host weighted by the K band luminosity as a proxy for the galaxies’ stellar masses. We also require the host to be at an appropriate distance and a similar declination to the original position of the trigger. The original skymap is then shifted across the sky to match the trigger’s new location (as the real position of the source is known relative to the skymap). By maintaining a similar declination, we minimise the effect of the shift on the GW localisation due to the locations of the GW detectors. This means we effectively only change the (fictional) time that a given trigger was detected which has no impact on the follow-up strategy. For sources that weren’t selected to be in a host galaxy, their position and associated skymap were left unchanged from the original simulation.
@@ 3.1 Kilonova component
Their tiny observed population mean that kilonovae remain poorly understood and models of their emission are diverse in behaviour (e.g. Rossi et al., 2022; Troja, 2023). For instance, toy models with assumed ejecta velocity distributions and opacities can approximate much of the behaviour of a kilonova (Metzger, 2019). More advanced radiation transfer models, on the other hand, can fully predict the angular dependence of luminosity and spectra (Kasen et al., 2017; Bulla, 2019; Korobkin et al., 2021; Wollaeger et al., 2021; Bulla, 2023). Arguably the most successful models, however, combine inferences and predictions from both the GW and EM data (Coughlin et al., 2019; Dietrich et al., 2020; Breschi et al., 2021). Here we use the analytical BNS models of Nicholl et al. (2021, hereafter N21) which offer a distinct advantage over these options - they are designed to generate a light curve from only parameters inferred from a GW signal. These models are therefore ideal for use in our planning and during O4 itself.
@@ 5.2 Binary black hole mergers
To this point we have only considered the gravitational wave sources detectable by LVK that are expected to produce significant EM emission: BNS and NSBH mergers. However, the vast majority of sources detected by LVK are binary black hole (BBH) mergers. Rather than the kilonova and short GRB afterglow emission components examined here, the possible EM emission from a BBH merger is far more unconstrained. Due to the inherent nature of a BBH system, EM emission can only be produced if the merger occurs in a dense environment, for instance, the disk of an active galactic nucleus (e.g. Kelly et al., 2017; McKernan et al., 2019). No confirmed EM counterpart to a BBH merger has thus far been discovered with only a few examples of possible candidates (e.g. Connaughton et al., 2016; Graham et al., 2020, 2023; Cabrera et al., 2024, see also Greiner et al. 2016 and Connaughton et al. 2018). Due to the lack of constraints on these triggers’ emission, Swift follows a different strategy to that explored in this work. Instead Swift follows-up only the best localised events with 90% error areas $\leq 30$ deg2and using minimal length 80 s exposures. This ensures the area is covered rapidly and minimises disruption to Swift’s other activities. During O4, Swift has followed up several BBH mergers meeting the criterion above, with no likely EM counterpart candidates so far.
@@ Acknowledgements
RAJEF acknowledges support from the UK Space Agency and the European Union’s Horizon 2020 Programme under the AHEAD2020 project (grant agreement number 871158). ET acknowledges support by the European Research Council through the Consolidator grant BHianca (Grant agreement ID: 101002761).